OpenStudy (jazzyfa30):

determine the quadrant in which theta lies if sec theta <0 and tan theta >0

OpenStudy (jazzyfa30):

@johnweldon1993

OpenStudy (johnweldon1993):

Well we know that \[\large sec(x) = \frac{1}{cos(x)}\] And we want that to be less than 0...so when is \[\large \frac{1}{cos(x)} < 0\]?

OpenStudy (jazzyfa30):

ok

OpenStudy (johnweldon1993):

 so so far we have quadrants 2 and 3

OpenStudy (johnweldon1993):

Now: \[\large tan(x) = \frac{sin(x)}{cos(x)}\] When is that greater than 0? well it would mean either BOTH cos and sin are negative..or they are BOTH positive... That happens